OpenStudy (anonymous):

Find the area of a rectangular object which has a length of 4√2 inches and a width of 6√6 inches. Explain, in complete sentences, how you arrived at the simplified answer and give the final solution in simplified radical form.

OpenStudy (anonymous):

@AriPotta

OpenStudy (espex):

The area of a rectangle is L x W, so take your 4√2 and multiply it by your 6√6, how far have you gotten in your solution?

OpenStudy (anonymous):

24√12***

OpenStudy (espex):

Yes, but now you can simplify the expression under the radical more.

OpenStudy (anonymous):

6√2

OpenStudy (espex):

Not quite. You have 24 √12, which means that you have 24 √(4*3)

OpenStudy (espex):

Which means that you have 24 * √4 * √3

OpenStudy (espex):

Take a look at the last response... What can you do with that?

OpenStudy (anonymous):

I really dont know

OpenStudy (espex):

Do you see how \[\sqrt{12}\] is just \[\sqrt{4*3}\]

OpenStudy (anonymous):

Yes

OpenStudy (espex):

And do you see how you can break them up like \[\sqrt{4}\times\sqrt{3}\]

OpenStudy (anonymous):

yes

OpenStudy (espex):

What is the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (espex):

Right, so now you have \[24\times 2 \times \sqrt{3}\]

OpenStudy (anonymous):

ok

OpenStudy (espex):

And your final answer is?

OpenStudy (anonymous):

48√3

OpenStudy (espex):

Precisely. Good job. :)